George is six times as old as Georgette. Two years ago, the sum of their ages was 31 years. How old was George 2 years ago. Please provide a chart in the beginning of the problem. THANKS!!!!!!
Answer by chibisan(131)

let george be 6x
so , georgette will be x
6x-2 + x-2 = 31
7x -4 = 31
7x = 35
x = 5
george 2years ago : (5*6)-2 = 28 years old.